### v1

# Investigating charmed hybrid baryons via QCD sum rules

Abstract
Motivated by recent interpretations of the $\eta_{1}(1855)$ and $X(2370)$ as a hybrid meson and a glueball, respectively, the possible existence of hybrid baryons has become a subject of considerable interest. In this letter, we investigate charmed hybrid baryons using the QCD sum rule method within the framework of heavy quark effective theory. We construct twenty-eight interpolating currents for charmed hybrid baryons, seven of which are employed in QCD sum rule analyses of nineteen states with quark-gluon configurations $qqcg$, $qscg$, and $sscg$ ($q=u/d$). The masses of the lowest-lying charmed hybrid baryons in the $SU(3)$ flavor $\mathbf{6}_{F}$ representation are calculated to be $M_{\Sigma_{cg}(1/2^{+})}=3.46^{+0.33}_{-0.21}~{}\rm{GeV}$, $M_{\Xi^{\prime}_{cg}(1/2^{+})}=3.60^{+0.48}_{-0.16}~{}\rm{GeV}$, and $M_{\Omega_{cg}(1/2^{+})}=3.77^{+0.31}_{-0.16}~{}\rm{GeV}$. We propose that future experiments search for these states via their $P$-wave decay channels $ND^{(*)}$, $\Lambda D^{(*)}$, and $\Xi D^{(*)}$, respectively. Such investigations would provide valuable insight into the role of gluonic excitations in hadron structure.
charmed hybrid baryon, QCD sum rules, heavy quark effective theory
Introduction — Quantum Chromodynamics (QCD) allows for the existence of color singlets, including conventional hadrons, multi-quark states, glueballs, and hybrids. Exploring exotic states beyond conventional hadrons is one of the most intriguing areas of research in hadronic physics, as such studies enhance our understanding of non-perturbative QCD. Gluonic degrees of freedom play a crucial role in these exotic states. The excitation of the gluonic field, analogous to the role of valence quarks, manifests in glueballs, hybrid mesons, and hybrid baryons. While substantial research has been conducted on hybrid mesons, relatively few studies have focused on hybrid baryons. Consequently, investigating hybrid baryons beyond conventional hadron configurations remains a significant and challenging topic in hadronic physics, requiring the identification of distinctive characteristics to confirm their existence. A common approach involves predicting the mass spectra and decay widths of hybrid baryons and offering specific recommendations for experimental searches.
In 2022, the BESIII collaboration observed the isoscalar state $\eta_{1}(1855)$ with the exotic quantum numbers $I^{G}J^{PC}=0^{+}1^{-+}$ BESIII:2022riz ; BESIII:2022iwi . Additionally, several collaborations IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi ; E852:1997gvf ; E852:1998mbq ; Aoyagi:1993kn ; CrystalBarrel:1998cfz have reported three isovector states—$\pi_{1}(1400)$, $\pi_{1}(1600)$, and $\pi_{1}(2015)$—all possessing the exotic quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$. These states are regarded as strong candidates for hybrid mesons. In 2024, the BESIII collaboration also reported the $X(2370)$ as a possible lightest pseudoscalar glueball BESIII:2023wfi ; Huang:2025pyv . These observations collectively support the potential existence of hybrid baryons, although no hybrid hadron has yet been conclusively identified in experiments. Ongoing efforts to search for hybrid baryons are underway at CLAS in Hall B of Jefferson Lab Lanza:2017qel ; Burkert:2018nvj , with the first results expected in the coming years. The hybrid hadron system has also attracted considerable interest from theorists. These exotic hadrons have been studied using various theoretical frameworks, including the bag model Kusaka:1990ha ; Barnes:1977hg ; Barnes:1982fj ; Barnes:1983xx ; Duck:1983ju ; Carlson:1983xr , lattice QCD Dudek:2012ag ; Bali:1993fb ; Chen:2005mg ; Morningstar:1999rf ; Gui:2019dtm ; Michael:1985ne ; Juge:2002br ; Lacock:1996ny ; Hedditch:2005zf ; Bernard:2003jd , the large-$N_{c}$ approach Chow:1998tq , the flux-tube model Page:2004gv ; Capstick:2002wm ; Capstick:1999qq ; Page:1998qt ; Isgur:1983wj ; Qiu:2022ktc , the hybrid quark model Khosonthongkee:2014gla ; Galata:2014typ ; Li:1991sh , the constituent gluon model Horn:1977rq ; Szczepaniak:2001rg ; Guo:2007sm , and QCD sum rules Zhao:2023imq ; Kisslinger:1995yw ; Kisslinger:2003hk ; Azizi:2017xyx ; Tan:2024grd ; Lian:2024fsb , among others. Additional discussions on related developments can be found in Refs. Cimino:2024bol ; Khan:2020ahz ; Kittel:2005jm ; Page:2002mt ; Barnes:2000vn ; Chen:2022asf , which provide insights into recent progress. Notably, most studies of hybrid baryons have concentrated on systems composed entirely of either light or heavy quarks. In contrast, our focus is on singly heavy hybrid baryons containing only one heavy quark. In such systems, the heavy quark effective theory (HQET) can be effectively applied, providing a promising framework for advancing our understanding of non-perturbative QCD.
In this letter, we systematically investigate charmed hybrid baryons using the QCD sum rule method within the framework of heavy quark effective theory. We construct twenty-eight interpolating currents for charmed hybrid baryons, seven of which are employed to perform sum rule analyses for nineteen states with quark-gluon configurations $qqcg$, $qscg$, and $sscg$ ($q=u/d$). We carry out comprehensive analyses of their mass spectra for quantum numbers $J^{P}=1/2^{+}$, $3/2^{+}$, and $5/2^{+}$, and determine their decay constants, which are essential for understanding their decay behaviors in future studies. In particular, we calculate the masses of the lowest-lying charmed hybrid baryons within the $SU(3)$ flavor $\mathbf{6}_{F}$ representation to be:
$M_{\Sigma_{cg}(1/2^{+})}$ $=$ $3.46^{+0.33}_{-0.21}~{}\rm{GeV}\,,$
$M_{\Xi^{\prime}_{cg}(1/2^{+})}$ $=$ $3.60^{+0.48}_{-0.16}~{}\rm{GeV}\,,$
$M_{\Omega_{cg}(1/2^{+})}$ $=$ $3.77^{+0.31}_{-0.16}~{}\rm{GeV}\,.$
All these states carry the quantum numbers $J^{P}=1/2^{+}$. We propose that future experimental studies search for them through their $P$-wave decay channels $ND^{(*)}$, $\Lambda D^{(*)}$, and $\Xi D^{(*)}$, respectively. Such investigations would be instrumental in deepening our understanding of gluonic field excitations within hadronic systems.
Categorization of charmed hybrid baryon currents — Heavy quark symmetry plays a crucial role in the study of heavy hadrons, and we apply it to systematically categorize the charmed hybrid baryon currents as follows. In the charmed hybrid baryon system, the heavy quark serves as a static color source, and the system’s properties are primarily determined by the light degrees of freedom.
Let us focus on the ground-state $Qq_{1}q_{2}g$ system, where $Q$ represents the heavy $charm$ quark, $q_{1}$ and $q_{2}$ represent the two valence light quarks ($up$, $down$, or $strange$), and $g$ represents the valence gluon. In the heavy quark limit, the total angular momentum of the system reduces to
$J=s_{Q}\otimes s_{q_{1}}\otimes s_{q_{2}}\otimes s_{g}=s_{Q}\otimes j_{l}\,,$ (1)
where $s_{Q}$, $s_{q_{1}}$, $s_{q_{2}}$, and $s_{g}$ denote the spins of the heavy quark $Q$, the light quarks $q_{1}$ and $q_{2}$, and the gluon $g$, respectively. Additionally, $j_{l}$ denotes the total angular momentum of the light degrees of freedom, expressed as
$j_{l}=s_{q_{1}}\otimes s_{q_{2}}\otimes s_{g}=j_{qq}\otimes s_{g}\,.$ (2)
Here, $j_{qq}=s_{q_{1}}\otimes s_{q_{2}}$ denotes the combined spin of the two light quarks. These two light quarks satisfy the exchange antisymmetry requirement for identical particles:
  • The flavor structure of the two light quarks can be either antisymmetric or symmetric. Accordingly, we use $\Lambda_{cg}$ and $\Xi_{cg}$ to denote the charmed hybrid baryons in the $SU(3)$ flavor $\bar{\textbf{3}}_{F}$ representation, while $\Sigma_{cg}$, $\Xi^{\prime}_{cg}$, and $\Omega_{cg}$ denote those in the $SU(3)$ flavor ${\textbf{6}}_{F}$ representation.
  • The color structure of the two light quarks can be either antisymmetric $(\bar{\textbf{3}}_{C})$ or symmetric $(\textbf{6}_{C})$, while the overall color structure of the three quarks is always $\textbf{8}_{C}$.
  • The spin structure of the two light quarks can be either antisymmetric ($j_{qq}=0$) or symmetric ($j_{qq}=1$).
Figure 1: Categorization of charmed hybrid baryon currents.
The charmed hybrid baryon currents can generally be constructed using one charm quark field $h_{v}^{a}(x)$, two light quark fields $q_{a}(x)$ and $q_{b}(x)$, and one gluon field strength tensor $G_{\mu\nu}^{n}(x)$ (or one dual gluon field strength tensor $\widetilde{G}_{\mu\nu}^{n}(x)=G^{n,\rho\sigma}(x)\times\epsilon_{\mu\nu\rho\sigma}/2$), with the color indices $a/b=1\cdots 3$ and $n=1\cdots 8$ as well as the Lorentz indices $\mu/\nu=0\cdots 3$. As illustrated in Fig. 1, we identify a total of twenty-eight charmed hybrid baryon currents, denoted as $J^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}/\widetilde{J}^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}$, which can be categorized into eight multiplets, denoted as $[F,C,j_{qq},j_{l},G/\widetilde{G}]$. Here, $F$ and $C$ represent the flavor and color representations of the two light quarks, while $J(\widetilde{J})$ and $G(\widetilde{G})$ indicate the presence of a gluon (or dual gluon) in the current. Their complete forms are provided in the supplemental material. We have systematically studied all these charmed hybrid baryon currents and found that the current $J_{6_{f},\bar{3}_{c},1,1,1/2^{+}}$, belonging to the doublet $[6_{f},\bar{3}_{c},1,1,G]$, couples to the lowest-lying charmed hybrid baryon. The explicit forms of this current and its spin partner for the $\Sigma_{cg}$ states are presented as follows ($q=up/down$):
$J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}=\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}]\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$ (3)
$J_{\Sigma_{cg},\bar{3}_{c},1,1,3/2^{+}}^{\alpha}=\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}](g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,.$
A point of confusion arises because these two currents exhibit a symmetric flavor structure for the two light quarks, which contrasts with the flavor structure of the lowest-lying charmed baryon $\Lambda_{c}$, but is consistent with the flavor structure of the $\Sigma_{c}$.
QCD sum rule analyses — We have systematically studied the charmed hybrid baryon currents given in Eqs. (3-4) and Eqs. (16-43) using the QCD sum rule method within the framework of heavy quark effective theory. As an example, we examine the current $J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}$ given in Eq. (3) as follows, which belongs to the doublet $[\Sigma_{cg},\bar{3}_{c},1,1,G]$. It couples to both the charmed hybrid baryon with $J^{P}=1/2^{+}$ and the one with $J^{P}=1/2^{-}$ through
$\langle 0|J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}|\Sigma_{cg}^{+}\rangle$ $=$ $f_{+}u\,,$ (5)
$\langle 0|J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}|\Sigma_{cg}^{-}\rangle$ $=$ $f_{-}\gamma_{5}u\,,$ (6)
where $f_{\pm}$ are the decay constants and $u$ is the Dirac spinor. Consequently, the two-point correlation function can be constructed as
$\Pi_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}(\omega)$ (7)
$=$ $i\int d^{4}xe^{ikx}\langle 0|T[J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}(x)\bar{J}_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}(0)]|0\rangle$
$=$ ${1+v\!\!\!/\over 2}\Pi_{+}(\omega)+{1-v\!\!\!/\over 2}\Pi_{-}(\omega)\,,$
where $\omega=v\cdot k$ is the external off-shell energy, and $\Pi_{\pm}$ are respectively contributed by $\Sigma_{cg}^{\pm}$. We calculate Eq. (7) at the quark-gluon level using the operator product expansion (OPE) method to derive
$\Pi_{-}(\omega)=0\,,$ (8)
suggesting that the current $J_{\Sigma_{cg},\bar{3}_{c},1,1/2^{+}}$ fully couples to the positive-parity state $\Sigma_{cg}^{+}$. This allows us to simplify Eq. (7) at the hadron level as
$\Pi_{+}(\omega)={f^{2}_{+}\over\bar{\Lambda}_{+}-\omega}+\rm{higher~{}~{}states}\,,$ (9)
with $\bar{\Lambda}_{+}$ the sum rule result at the leading order. After performing the Borel transformation at both the hadron and quark-gluon levels, we obtain
$\Pi_{+}(\omega_{c},T)$ $=$ $f_{+}^{2}e^{-\bar{\Lambda}_{+}/T}$ (10)
$=$ $\int_{0}^{\omega_{c}}e^{-\omega/T}d\omega\times\Big{(}{\alpha_{s}\over 1890\pi^{5}}\omega^{9}+{\langle g_{s}^{2}GG\rangle\over 240\pi^{4}}\omega^{9}$
$-{\alpha_{s}\langle g_{s}^{2}GG\rangle\over 1440\pi^{5}}\omega^{5}-{\langle g_{s}^{3}G^{3}\rangle\over 192\pi^{4}}\omega^{3}+{8\alpha_{s}\langle\bar{q}q\rangle^{2}\over 9\pi}\omega^{3}$
$+{\langle g_{s}^{2}GG\rangle^{2}\over 4608\pi^{4}}\omega-{41\alpha_{s}\langle\bar{q}q\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 48\pi}\omega\Big{)}$
$+{\langle\bar{q}q\rangle^{2}\langle g_{s}^{2}GG\rangle\over 72}-{\langle\bar{q}q\rangle\langle g_{s}^{2}GG\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 576T^{2}}$
$-{\langle\bar{q}q\rangle^{2}\langle g_{s}^{3}G^{3}\rangle\over 576T^{2}}+{25\alpha_{s}\langle\bar{q}\sigma Gq\rangle^{2}\over 768\pi}\,,$
from which we further derive
$\overline{\Lambda}_{+}(\omega_{c},T)$ $=$ $\frac{1}{\Pi_{+}(\omega_{c},T)}\times\frac{\partial\Pi_{+}(\omega_{c},T)}{\partial(-1/T)}\,,$ (11)
$f_{+}^{2}(\omega_{c},T)$ $=$ $\Pi_{+}(\omega_{c},T)\times e^{\overline{\Lambda}_{+}(\omega_{c},T)/T}\,.$ (12)
In the above expressions, we have evaluated the two-point correlation function up to the twelfth dimension, including the quark condensate $\langle\bar{q}q\rangle$, the two-gluon condensate $\langle g_{s}^{2}GG\rangle$, the three-gluon condensate $\langle g_{s}^{3}G^{3}\rangle$, the quark-gluon mixed condensate $\langle g_{s}\bar{q}\sigma Gq\rangle$, and their combinations.
Additionally, we have evaluated the sum rule result at the $\mathcal{O}(1/m_{c})$ order, and the mass of the $\Sigma_{cg}^{+}$ state is given by
$M_{+}(\omega_{c},T)=m_{c}+\bar{\Lambda}_{+}(\omega_{c},T)+\delta M_{+}(\omega_{c},T)\,.$ (13)
There are two free parameters, the threshold value $\omega_{c}$ and the Borel mass $T$, and we need to find their proper working regions. Detailed analyses are provided in the supplemental material, and the numerical results for the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$ are provided in Table 1.
Table 1: Parameters of the charmed hybrid baryons belonging to the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$, calculated using the QCD sum rule method within the framework of heavy quark effective theory.
   Multiplets     $\omega_{c}$     Working region         $\overline{\Lambda}$       State     Mass     Difference    Decay constant
(GeV) (GeV) (GeV) (GeV) (GeV) (GeV4)
$[\bar{3}_{f},\bar{3}_{c},0,1,G]$ 3.20 $0.400\leq T\leq 0.410$ $2.72^{+0.17}_{-0.15}$ $\Lambda_{cg}(1/2^{+})$ $4.28^{+0.21}_{-0.19}$ $0.31^{+0.06}_{-0.05}$ $0.155^{+0.031}_{-0.024}$
$\Lambda_{cg}(3/2^{+})$ $4.59^{+0.24}_{-0.20}$ $0.045^{+0.009}_{-0.007}$
3.45 $0.409\leq T\leq 0.437$ $2.76^{+0.11}_{-0.09}$ $\Xi_{cg}(1/2^{+})$ $4.38^{+0.14}_{-0.12}$ $0.27^{+0.05}_{-0.04}$ $0.176^{+0.023}_{-0.019}$
$\Xi_{cg}(3/2^{+})$ $4.65^{+0.15}_{-0.13}$ $0.051^{+0.007}_{-0.005}$
$[6_{f},\bar{3}_{c},1,0,G]$ $2.80$ $0.394\leq T\leq 0.401$ $2.25\pm 0.11$ $\Sigma_{cg}(1/2^{+})$ $4.06^{+0.16}_{-0.15}$ $0.093^{+0.014}_{-0.012}$
3.05 $0.373\leq T\leq 0.425$ $2.32^{+0.14}_{-0.06}$ $\Xi^{\prime}_{cg}(1/2^{+})$ $4.17^{+0.16}_{-0.10}$ $0.109^{+0.019}_{-0.011}$
3.30 $0.357\leq T\leq 0.447$ $2.44^{+0.15}_{-0.08}$ $\Omega_{cg}(1/2^{+})$ $4.34^{+0.17}_{-0.11}$ $0.191^{+0.041}_{-0.020}$
$[6_{f},\bar{3}_{c},1,1,G]$ 2.95 $0.368\leq T\leq 0.398$ $2.46^{+0.30}_{-0.14}$ $\Sigma_{cg}(1/2^{+})$ $3.46^{+0.33}_{-0.21}$ $1.07^{+0.17}_{-0.14}$ $0.079^{+0.033}_{-0.030}$
$\Sigma_{cg}(3/2^{+})$ $4.54^{+0.38}_{-0.20}$ $0.046^{+0.019}_{-0.018}$
3.20 $0.351\leq T\leq 0.423$ $2.50^{+0.45}_{-0.08}$ $\Xi^{\prime}_{cg}(1/2^{+})$ $3.60^{+0.48}_{-0.16}$ $1.07^{+0.19}_{-0.13}$ $0.090^{+0.073}_{-0.010}$
$\Xi^{\prime}_{cg}(3/2^{+})$ $4.67^{+0.57}_{-0.12}$ $0.052^{+0.042}_{-0.006}$
3.45 $0.351\leq T\leq 0.447$ $2.62^{+0.28}_{-0.08}$ $\Omega_{cg}(1/2^{+})$ $3.77^{+0.31}_{-0.16}$ $1.07^{+0.16}_{-0.13}$ $0.158^{+0.072}_{-0.016}$
$\Omega_{cg}(3/2^{+})$ $4.84^{+0.35}_{-0.12}$ $0.091^{+0.042}_{-0.009}$
$[6_{f},\bar{3}_{c},1,2,G]$ 3.50 $0.418\leq T\leq 0.429$ $2.98^{+0.16}_{-0.14}$ $\Sigma_{cg}(3/2^{+})$ $5.36^{+0.24}_{-0.20}$ $-0.70^{+0.11}_{-0.11}$ $0.364^{+0.063}_{-0.051}$
$\Sigma_{cg}(5/2^{+})$ $4.66^{+0.21}_{-0.18}$ $0.123^{+0.021}_{-0.017}$
3.75 $0.428\leq T\leq 0.457$ $3.04^{+0.10}_{-0.09}$ $\Xi^{\prime}_{cg}(3/2^{+})$ $5.45^{+0.16}_{-0.14}$ $-0.67^{+0.09}_{-0.10}$ $0.422^{+0.052}_{-0.043}$
$\Xi^{\prime}_{cg}(5/2^{+})$ $4.78^{+0.15}_{-0.13}$ $0.143^{+0.018}_{-0.015}$
4.00 $0.434\leq T\leq 0.484$ $3.16\pm 0.09$ $\Omega_{cg}(3/2^{+})$ $5.63^{+0.15}_{-0.14}$ $-0.65^{+0.09}_{-0.10}$ $0.740^{+0.080}_{-0.073}$
$\Omega_{cg}(5/2^{+})$ $4.98^{+0.14}_{-0.13}$ $0.250^{+0.027}_{-0.025}$
Summary and discussions — In recent years, substantial theoretical and experimental efforts have been devoted to the study of glueballs and hybrid mesons. However, their internal structures remain elusive, underscoring the necessity for continued investigations—particularly into the comparatively less explored hybrid baryons. In this work, we employ the QCD sum rule method within the framework of heavy quark effective theory to investigate charmed hybrid baryons with various quantum numbers. We systematically construct twenty-eight interpolating currents for these baryons and utilize seven of them to compute the masses of nineteen states with quark-gluon configurations $qqcg$, $qscg$, and $sscg$ ($q=u/d$). The resulting mass predictions are summarized in Table 1.
The charmed hybrid baryons belonging to the doublet $[6_{f},\bar{3}c,1,1,G]$ comprise six states: $\Sigma_{cg}(1/2^{+})$, $\Sigma_{cg}(3/2^{+})$, $\Xi^{\prime}_{cg}(1/2^{+})$, $\Xi^{\prime}_{cg}(3/2^{+})$, $\Omega_{cg}(1/2^{+})$, and $\Omega_{cg}(3/2^{+})$. Among these, the three spin-$1/2$ states are predicted to be the lowest in mass:
$M_{\Sigma_{cg}(1/2^{+})}$ $=$ $3.46^{+0.33}_{-0.21}~{}\rm{GeV}\,,$
$M_{\Xi^{\prime}_{cg}(1/2^{+})}$ $=$ $3.60^{+0.48}_{-0.16}~{}\rm{GeV}\,,$
$M_{\Omega_{cg}(1/2^{+})}$ $=$ $3.77^{+0.31}_{-0.16}~{}\rm{GeV}\,.$
These hybrid baryons can decay via the excitation of a color-octet $\bar{q}q$ or $\bar{s}s$ pair from the valence gluon, followed by the recombination of the remaining color-octet clusters—$qqc$, $qsc$, or $ssc$—with the excited $\bar{q}q$ or $\bar{s}s$ pair into color-singlet baryons and mesons. Representative decay channels for these states are listed in Table 2, offering viable avenues for their experimental identification in future studies.
Table 2: Possible decay channels of the charmed hybrid baryons belonging to the doublet $[6_{f},\bar{3}_{c},1,1,G]$.
States Decay patterns
$\Sigma_{cg}(1/2^{+})$ $ND$, $\Sigma D_{s}$, $ND^{*}$, $\Sigma D_{s}^{*}$, $\Delta D$, $\Sigma^{*}D_{s}$
$\Xi^{\prime}_{cg}(1/2^{+})$ $\Lambda D$, $\Xi D_{s}$, $\Lambda D^{*}$, $\Xi D_{s}^{*}$, $\Xi^{*}D_{s}$
$\Omega_{cg}(1/2^{+})$ $\Xi D$, $\Xi D^{*}$, $\Xi^{*}D$, $\Omega D_{s}$, $\Xi^{*}D^{*}$
ACKNOWLEDGMENTS
We are grateful to Shi-Lin Zhu for the helpful discussions. This project is supported by the China Postdoctoral Science Foundation under Grants No. 2024M750049, the National Natural Science Foundation of China under Grant No. 12075019, and the Fundamental Research Funds for the Central Universities.
Supplemental Material
This supplemental material provides additional details on our framework. In this Letter we apply the QCD sum rule method to study the charmed hybrid baryons within the framework of heavy quark effective theory. Their interpolating currents can generally be constructed by using one charm quark field $h_{v}^{a}(x)$, two light quark fields $q_{a}(x)$ and $q_{b}(x)$, and one gluon field strength tensor $G_{\mu\nu}^{n}(x)$ (or one dual gluon field strength tensor $\widetilde{G}_{\mu\nu}^{n}(x)=G^{n,\rho\sigma}(x)\times\epsilon_{\mu\nu\rho\sigma}/2$), with the color indices $a/b=1\cdots 3$ and $n=1\cdots 8$ as well as the Lorentz indices $\mu/\nu=0\cdots 3$. The (dual) gluon field is a color octet, so the combination of the three quark fields should also form a color octet, with two possible combinations:
$\epsilon_{abe}[q^{a}q^{b}]{\lambda_{ec}^{n}\over 2}h_{v}^{c}~{}\times~{}G^{n}/\widetilde{G}^{n}\,,$ (14)
$\epsilon_{aec}[q^{a}q^{b}+q^{b}q^{a}]{\lambda_{eb}^{n}\over 2}h_{v}^{c}~{}\times~{}G^{n}/\widetilde{G}^{n}\,.$ (15)
The light diquark in the former combination is a color anti-triplet, while the one in the latter is a color sextet.
As summarized in Fig. 1, we identify a total of twenty-eight charmed hybrid baryon currents, denoted as $J^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}/\widetilde{J}^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}$, which can be categorized into eight multiplets, denoted as $[F,C,j_{qq},j_{l},G/\widetilde{G}]$. The explicit expressions of these currents are given as follows:
  • The doublet $[\bar{3}_{f},\bar{3}_{c},0,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[\bar{3}_{f},\bar{3}_{c},0,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $J_{\bar{3}_{f},\bar{3}_{c},0,1,1/2^{+}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{5}q^{b}]\sigma_{t}^{\mu\nu}g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,,$ (16)
    $\widetilde{J}_{\bar{3}_{f},\bar{3}_{c},0,1,1/2^{-}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{5}q^{b}]\sigma_{t}^{\mu\nu}g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,,$ (17)
    $J_{\bar{3}_{f},\bar{3}_{c},0,1,3/2^{+}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{5}q^{b}]$ (18)
    $\times$ $(g_{t}^{\alpha\mu}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\mu})\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,,$
    $\widetilde{J}_{\bar{3}_{f},\bar{3}_{c},0,1,3/2^{-}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{5}q^{b}]$ (19)
    $\times$ $(g_{t}^{\alpha\mu}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\mu})\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,.$
  • The singlet $[6_{f},\bar{3}_{c},1,0,G]$ contains one state with $J^{P}=1/2^{+}$, and its dual singlet $[6_{f},\bar{3}_{c},1,0,\widetilde{G}]$ contains another state with $J^{P}=1/2^{-}$. Their corresponding currents are
    $J_{6_{f},\bar{3}_{c},1,0,1/2^{+}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\sigma_{t}^{\mu\nu}\gamma_{5}q^{b}]g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,,$ (20)
    $\widetilde{J}_{6_{f},\bar{3}_{c},1,0,1/2^{-}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\sigma_{t}^{\mu\nu}\gamma_{5}q^{b}]g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,.$ (21)
  • The doublet $[6_{f},\bar{3}_{c},1,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[6_{f},\bar{3}_{c},1,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $J_{6_{f},\bar{3}_{c},1,1,1/2^{+}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}]\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,,$ (22)
    $\widetilde{J}_{6_{f},\bar{3}_{c},1,1,1/2^{-}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}]\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,,$ (23)
    $J_{6_{f},\bar{3}_{c},1,1,3/2^{+}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}]$ (24)
    ×$~{}~{}~{}~{}\times~{}~{}(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,,$
    $\widetilde{J}_{6_{f},\bar{3}_{c},1,1,3/2^{-}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}]$ (25)
    ×$~{}~{}~{}~{}\times~{}~{}(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,.$
  • The doublet $[6_{f},\bar{3}_{c},1,2,G]$ contains two states with $J^{P}=3/2^{+}$ and $5/2^{+}$. Its dual doublet $[6_{f},\bar{3}_{c},1,2,\widetilde{G}]$ contains two other states with $J^{P}=3/2^{-}$ and $5/2^{-}$. Their corresponding currents are
    $J_{6_{f},\bar{3}_{c},1,2,3/2^{+}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}]$ (26)
    $\times$ $\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{n}\over 2}G_{\nu_{1}\nu_{2}}^{n}h_{v}^{c}\,,$
    $\widetilde{J}_{6_{f},\bar{3}_{c},1,2,3/2^{-}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}]$ (27)
    $\times$ $\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{n}h_{v}^{c}\,,$
    $J_{6_{f},\bar{3}_{c},1,2,5/2^{+}}^{\alpha\beta}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}]$ (28)
    $\times$ $\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{n}\over 2}G_{\nu_{1}\nu_{2}}^{n}h_{v}^{c}\,,$
    $\widetilde{J}_{6_{f},\bar{3}_{c},1,2,5/2^{-}}^{\alpha\beta}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}]$ (29)
    $\times$ $\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{n}h_{v}^{c}\,.$
  • The doublet $[6_{f},6_{c},0,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[6_{f},6_{c},0,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $J_{6_{f},6_{c},0,1,1/2^{+}}$ $=$ $\epsilon_{aec}[q^{aT}\mathcal{C}\gamma_{5}q^{b}+q^{bT}\mathcal{C}\gamma_{5}q^{a}]$ (30)
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\sigma_{t}^{\mu\nu}g_{s}{\lambda_{eb}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{6_{f},6_{c},0,1,1/2^{-}}$ $=$ $\epsilon_{aec}[q^{aT}\mathcal{C}\gamma_{5}q^{b}+q^{bT}\mathcal{C}\gamma_{5}q^{a}]$ (31)
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\sigma_{t}^{\mu\nu}g_{s}{\lambda_{eb}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,,$
    $J_{6_{f},6_{c},0,1,3/2^{+}}^{\alpha}$ $=$ $\epsilon_{aec}[q^{aT}\mathcal{C}\gamma_{5}q^{b}+q^{bT}\mathcal{C}\gamma_{5}q^{a}]$ (32)
    $\times$ $(g_{t}^{\alpha\mu}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\mu})\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{eb}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{6_{f},6_{c},0,1,3/2^{-}}^{\alpha}$ $=$ $\epsilon_{aec}[q^{aT}\mathcal{C}\gamma_{5}q^{b}+q^{bT}\mathcal{C}\gamma_{5}q^{a}]$ (33)
    $\times$ $(g_{t}^{\alpha\mu}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\mu})\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{eb}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,.$
  • The singlet $[\bar{3}_{f},6_{c},1,0,G]$ contains one state with $J^{P}=1/2^{+}$, and its dual singlet $[\bar{3}_{f},6_{c},1,0,\widetilde{G}]$ contains another state with $J^{P}=1/2^{-}$. Their corresponding currents are
    $J_{\bar{3}_{f},6_{c},1,0,1/2^{+}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\sigma_{t}^{\mu\nu}\gamma_{5}q^{b}+q^{bT}\mathcal{C}\sigma_{t}^{\mu\nu}\gamma_{5}q^{a}]$ (34)
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{\bar{3}_{f},6_{c},1,0,1/2^{-}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\sigma_{t}^{\mu\nu}\gamma_{5}q^{b}+q^{bT}\mathcal{C}\sigma_{t}^{\mu\nu}\gamma_{5}q^{a}]$ (35)
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,.$
  • The doublet $[\bar{3}_{f},6_{c},1,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[\bar{3}_{f},6_{c},1,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    $J_{\bar{3}_{f},6_{c},1,1,1/2^{+}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}+q^{bT}\mathcal{C}\gamma_{t}^{\mu}q^{a}]$ (36)
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{\bar{3}_{f},6_{c},1,1,1/2^{-}}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}+q^{bT}\mathcal{C}\gamma_{t}^{\mu}q^{a}]$ (37)
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,,$
    $J_{\bar{3}_{f},6_{c},1,1,3/2^{+}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}+q^{bT}\mathcal{C}\gamma_{t}^{\mu}q^{a}]$ (38)
    $\times$ $(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{\bar{3}_{f},6_{c},1,1,3/2^{-}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t}^{\mu}q^{b}+q^{bT}\mathcal{C}\gamma_{t}^{\mu}q^{a}]$ (39)
    $\times$ $(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,.$
  • The doublet $[\bar{3}_{f},6_{c},1,2,G]$ contains two states with $J^{P}=3/2^{+}$ and $5/2^{+}$. Its dual doublet $[\bar{3}_{f},6_{c},1,2,\widetilde{G}]$ contains two other states with $J^{P}=3/2^{-}$ and $5/2^{-}$. Their corresponding currents are
    $J_{\bar{3}_{f},6_{c},1,2,3/2^{+}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}+q^{bT}\mathcal{C}\gamma_{t,\mu_{1}}q^{a}]$ (40)
    ×$~{}\times~{}~{}\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{m}\over 2}G_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{\bar{3}_{f},6_{c},1,2,3/2^{-}}^{\alpha}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}+q^{bT}\mathcal{C}\gamma_{t,\mu_{1}}q^{a}]$ (41)
    ×$~{}\times~{}~{}\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,,$
    $J_{\bar{3}_{f},6_{c},1,2,5/2+}^{\alpha\beta}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}+q^{bT}\mathcal{C}\gamma_{t,\mu_{1}}q^{a}]$ (42)
    ×$~{}~{}~{}~{}\times~{}~{}\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,,$
    $\widetilde{J}_{\bar{3}_{f},6_{c},1,2,5/2-}^{\alpha\beta}$ $=$ $\epsilon_{abe}[q^{aT}\mathcal{C}\gamma_{t,\mu_{1}}q^{b}+q^{bT}\mathcal{C}\gamma_{t,\mu_{1}}q^{a}]$ (43)
    ×$~{}~{}~{}~{}\times~{}~{}\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,.$
In the above expressions, $\gamma_{t}^{\mu}=\gamma^{\mu}-v\!\!\!/v^{\mu}$, $g_{t}^{\mu\nu}=g^{\mu\nu}-v^{\mu}v^{\nu}$, $\sigma_{t}^{\mu\nu}=\sigma^{\mu\nu}-\sigma^{\mu\alpha}v_{\alpha}v^{\nu}-\sigma^{\alpha\nu}v_{\alpha}v^{\mu}$, and $G_{\mu\nu}^{n}=\partial_{\mu}A_{\nu}^{n}-\partial_{\nu}A_{\mu}^{n}+g_{s}f^{npq}A_{p,\mu}A_{q,\nu}$; $\Gamma_{J=2}^{\alpha\beta,\mu\nu}$ and $\Gamma_{J=5/2}^{\alpha\beta,\mu\nu}$ are the $J=2$ and $J=5/2$ projection operators:
$\Gamma_{J=2}^{\alpha\beta,\mu\nu}$ $=$ $g_{t}^{\alpha\mu}g_{t}^{\beta\nu}+g_{t}^{\alpha\nu}g_{t}^{\beta\mu}-{2\over 3}g_{t}^{\alpha\beta}g_{t}^{\mu\nu}\,,$ (44)
$\Gamma_{J=5/2}^{\alpha\beta,\mu\nu}$ $=$ $g_{t}^{\alpha\mu}g_{t}^{\beta\nu}+g_{t}^{\alpha\nu}g_{t}^{\beta\mu}-{2\over 5}g_{t}^{\alpha\beta}g_{t}^{\mu\nu}$
-- ${1\over 5}g_{t}^{\alpha\mu}\gamma_{t}^{\beta}\gamma_{t}^{\nu}-{1\over 5}g_{t}^{\alpha\nu}\gamma_{t}^{\beta}\gamma_{t}^{\mu}-{1\over 5}g_{t}^{\beta\mu}\gamma_{t}^{\alpha}\gamma_{t}^{\nu}-{1\over 5}g_{t}^{\beta\nu}\gamma_{t}^{\alpha}\gamma_{t}^{\mu}\,.$
In this Letter we apply the QCD sum rule method to systematically study the charmed hybrid baryon currents listed above within the framework of heavy quark effective theory. In particular, the numerical results for the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$ are presented in Table 1. As discussed previously, these currents can couple to charmed hybrid baryons with both positive and negative parities. For example, the positive-parity current $J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}$ generally couples to both the positive-parity state $B_{+}$ and the negative-parity state $B_{-}$ through
$\langle 0|J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}|B_{+}\rangle$ $=$ $f^{+}_{F,C,j_{qq},j_{l},J^{+}}u^{\alpha_{1}\cdots\alpha_{J-1/2}}\,,$ (46)
$\langle 0|J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}|B_{-}\rangle$ $=$ $f^{-}_{F,C,j_{qq},j_{l},J^{+}}\gamma_{5}u^{\alpha_{1}\cdots\alpha_{J-1/2}}\,.$ (47)
Its two-point correlation function can generally be written as
$\Pi_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2},\beta_{1},\cdots\beta_{J-1/2}}(\omega)$ (48)
$=$ $i\int d^{4}xe^{ikx}\langle 0|T[J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}(x)\bar{J}_{F,C,j_{qq},j_{l},J^{+}}^{\beta_{1},\cdots\beta_{J-1/2}}(0)]|0\rangle$
$=$ $\mathbb{S}[g_{t}^{\alpha_{1}\beta_{1}}\cdots g_{t}^{\alpha_{J-1/2}\beta_{J-1/2}}]$
$\times$ $\Big{(}{1+v\!\!\!/\over 2}\Pi^{+}_{F,C,j_{qq},j_{l},J^{+}}(\omega)+{1-v\!\!\!/\over 2}\Pi^{-}_{F,C,j_{qq},j_{l},J^{+}}(\omega)\Big{)}\,,$
where $\mathbb{S}[\cdots]$ denotes the symmetrization and the subtraction of trace terms in the sets ($\alpha_{1}\cdots\alpha_{J-1/2}$) and ($\beta_{1}\cdots\beta_{-1/2}$).
At the quark-gluon level, we calculate Eq. (48) using the operator product expansion (OPE) method, and the sum rule equations for the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$ are provided in the supplemental Mathematica file “OPE.nb”. During the calculations, we evaluate the OPE at the leading order, and the sum rule equation for the current $J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}$ is given in Eq. (10); we also evaluate the OPE at the $\mathcal{O}(1/m_{c})$ order, and the sum rule equations for this current are
$\delta_{\mathcal{K}}\Pi_{+}(\omega_{c},T)=f_{+}^{2}K_{+}e^{-\bar{\Lambda}_{+}/T}$ (50)
$=$ $\int_{0}^{\omega_{c}}e^{-\omega/T}d\omega\times\Big{(}-{\alpha_{s}\over 4950\pi^{5}}\omega^{11}-{11\langle g_{s}^{2}GG\rangle\over 5040\pi^{4}}\omega^{7}$
$+{13\alpha_{s}\langle g_{s}^{2}GG\rangle\over 24192\pi^{5}}\omega^{7}+{\langle g_{s}^{3}G^{3}\rangle\over 1920\pi^{4}}\omega^{5}-{8\alpha_{s}\langle\bar{q}q\rangle^{2}\over 9\pi}\omega^{5}$
$+{\langle g_{s}^{2}GG\rangle^{2}\over 256\pi^{4}}\omega^{3}-{47\langle g_{s}^{2}GG\rangle^{2}\over 110592\pi^{4}}\omega^{3}$
$+{269\alpha_{s}\langle\bar{q}q\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 144\pi}\omega^{3}-{41\alpha_{s}\langle g_{s}\bar{q}\sigma Gq\rangle^{2}\over 192\pi}\omega\Big{)}$
$-{5\langle\bar{q}q\rangle\langle g_{s}^{2}GG\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 512}-{\langle\bar{q}q\rangle^{2}\langle g_{s}^{3}G^{3}\rangle\over 288}\,,$
$\delta_{\mathcal{S}}\Pi_{+}(\omega_{c},T)=4\times f_{+}^{2}\Sigma_{+}e^{-\bar{\Lambda}_{+}/T}$
$=$ $\int_{0}^{\omega_{c}}e^{-\omega/T}d\omega\times\Big{(}{815\alpha_{s}\langle g_{s}^{2}GG\rangle\over 24192\pi^{5}}\omega^{7}+{\langle g_{s}^{3}G^{3}\rangle\over 960\pi^{4}}\omega^{5}$
$+{251\langle g_{s}^{2}GG\rangle^{2}\over 221184\pi^{4}}\omega^{3}-{511\alpha_{s}\langle\bar{q}q\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 108\pi}\omega^{3}$
$+{8779\alpha_{s}\langle g_{s}\bar{q}\sigma Gq\rangle^{2}\over 4608\pi}\omega+{9\alpha_{s}\langle\bar{q}q^{2}\rangle\langle g_{s}^{2}GG\rangle\over 32\pi}\omega\Big{)}$
$-{31\langle\bar{q}q\rangle\langle g_{s}^{2}GG\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 6912}+{\langle\bar{q}q\rangle^{2}\langle g_{s}^{3}G^{3}\rangle\over 288}$
$-{9\alpha_{s}\langle\bar{q}q\rangle\langle g_{s}^{2}GG\rangle\langle g_{s}\bar{q}\sigma Gq\rangle\over 256\pi}\,.$
Note that in the above expressions, we have considered only the $\Sigma_{cg}^{+}$ state with the positive parity and neglected the $\Sigma_{cg}^{-}$ state with the negative parity.
To calculate Eq. (50) and Eq. (50), we consider the following Lagrangian from the heavy quark effective theory
$\mathcal{L}_{\rm eff}=\bar{h}_{v}^{c}iv\cdot D_{t}h_{v}^{c}+\frac{1}{2m_{c}}\mathcal{K}+\frac{1}{2m_{c}}\mathcal{S}\,,$ (51)
where $\mathcal{K}$ represents the operator of nonrelativistic kinetic energy and $\mathcal{S}$ represents the Pauli term describing the chromomagnetic interaction:
$\mathcal{K}$ $=$ $\bar{h}_{v}^{c}(iD_{t})^{2}h_{v}^{c}\,,$ (52)
$\mathcal{S}$ $=$ $\frac{g_{s}}{2}C_{mag}(m_{c}/\mu)\bar{h}_{v}^{c}\sigma_{\mu\nu}G^{\mu\nu}h_{v}^{c}\,,$ (53)
with $C_{mag}(m_{c}/\mu)=[\alpha_{s}(m_{c})/\alpha_{s}(\mu)]^{3/\beta_{0}}$ and $\beta_{0}=11-2n_{f}/3$.
Following Eq. (9), we can express the pole term up to the ${\mathcal{O}}(1/m_{c})$ order as
$\Pi_{+}(\omega)$ $=$ $\frac{(f_{+}+\delta f_{+})^{2}}{\overline{\Lambda}_{+}+\delta M_{+}-\omega}$ (54)
$=$ $\frac{f_{+}^{2}}{\overline{\Lambda}_{+}-\omega}-\frac{\delta M_{+}f_{+}^{2}}{(\overline{\Lambda}_{+}-\omega)^{2}}+\frac{f_{+}\delta f_{+}}{\overline{\Lambda}_{+}-\omega}\,,$
where $\delta M_{+}$ and $\delta f_{+}$ denote the corrections to the mass and decay constant, respectively. To calculate $\delta M_{+}$, we consider the following three-point correlation functions
$\delta_{\mathcal{O}}\Pi_{F,C,j_{qq},j_{l},J^{P}}^{\alpha_{1}\cdots\alpha_{J-1/2},\beta_{1}\cdots\beta_{J-1/2}}(\omega,\omega^{\prime})$ (55)
$=$ $i^{2}\int d^{4}xd^{4}ye^{ikx-ik^{\prime}y}\langle 0|TJ_{F,C,j_{qq},j_{l},J^{P}}^{\alpha_{1}\cdots\alpha_{J-1/2}}(x)\mathcal{O}\bar{J}_{F,C,j_{qq},j_{l},J^{P}}^{\beta_{1}\cdots\beta_{J-1/2}}(y)|0\rangle$
$=$ $\mathbb{S}[g_{t}^{\alpha_{1}\beta_{1}}\cdots g_{t}^{\alpha_{J-1/2}\beta_{J-1/2}}]$
$\times$ $\left({1+v\!\!\!/\over 2}\delta_{\mathcal{O}}\Pi^{+}_{F,C,j_{qq},j_{l},J^{P}}+{1-v\!\!\!/\over 2}\delta_{\mathcal{O}}\Pi^{-}_{F,C,j_{qq},j_{l},J^{P}}\right)\,,$
where $\mathcal{O}=\mathcal{K}$ or $\mathcal{S}$, $k^{\prime}=k+q$, $\omega=v\cdot k$, and $\omega^{\prime}=v\cdot k^{\prime}$.
Based on Eq. (51), we can express Eq. (55) at the hadron level as
$\delta_{\mathcal{K}}\Pi_{+}$ $=$ $\frac{f_{+}^{2}K_{+}}{(\overline{\Lambda}_{+}-\omega)(\overline{\Lambda}_{+}-\omega^{\prime})}+\frac{f_{+}^{2}G_{\mathcal{K}}}{\overline{\Lambda}_{+}-\omega}+\frac{f_{+}^{2}G^{\prime}_{\mathcal{K}}}{\overline{\Lambda}_{+}-\omega^{\prime}}\,,$ (56)
$\delta_{\mathcal{S}}\Pi_{+}$ $=$ $\frac{d_{M}f_{+}^{2}\Sigma_{+}}{(\overline{\Lambda}_{+}-\omega)(\overline{\Lambda}_{+}-\omega^{\prime})}+\frac{d_{M}f_{+}^{2}G_{\mathcal{S}}}{\overline{\Lambda}_{+}-\omega}\,+\frac{d_{M}f_{+}^{2}G^{\prime}_{\mathcal{S}}}{\overline{\Lambda}_{+}-\omega^{\prime}}\,,$
where $K_{+}$, $\Sigma_{+}$, and $d_{M}$ are defined as
$K_{+}$$\equiv$ $\langle\Sigma_{cg}^{+}|\bar{h}_{v}^{c}(iD_{t})^{2}h_{v}^{c}|\Sigma_{cg}^{+}\rangle\,,$
$d_{M}\Sigma_{+}$$\equiv$ $\langle\Sigma_{cg}^{+}|{g_{s}\over 2}\bar{h}_{v}^{c}\sigma_{\mu\nu}G^{\mu\nu}h_{v}^{c}|\Sigma_{cg}^{+}\rangle\,,$
$d_{M}$$\equiv$ $d_{j,j_{l}}\,,$ (58)
$d_{j_{l}-1/2,j_{l}}$ $=$ $2j_{l}+2\,,$
$d_{j_{l}+1/2,j_{l}}$ $=$ $-2j_{l}\,.$
Note that the term $\mathcal{S}$ causes a mass splitting within the same doublet, while it can also lead to a mixing of states with the same spin-parity quantum number. We have accounted for the former effect, but the latter effect is found in Ref. Dai:1998ve to be negligible and is therefore not considered in the present study.
After performing the double Borel transformation to convert $\omega$ and $\omega^{\prime}$ into $T_{1}$ and $T_{2}$, and setting these two Borel parameters equal, we obtain Eq. (50) and Eq. (50). These two equations represent the sum rule results at the $\mathcal{O}(1/m_{c})$ order, from which we derive
$\delta M_{+}(\omega_{c},T)=-\frac{1}{4m_{c}}\left(K_{+}(\omega_{c},T)+d_{M}C_{mag}\Sigma_{+}(\omega_{c},T)\right)\,.$ (59)
To perform numerical analyses, we use the following values for various quark and gluon condensate ParticleDataGroup:2020ssz ; Ovchinnikov:1988gk ; Yang:1993bp ; Ellis:1996xc ; Ioffe:2002be ; Jamin:2002ev ; Gimenez:2005nt ; Narison:2011xe ; Narison:2018dcr :
$\langle\bar{q}q\rangle$ $=$ $-(0.240\pm 0.010)^{3}~{}\rm{GeV}\,,$
$\langle\bar{s}s\rangle$ $=$ $(0.8\pm 0.1)\times\langle\bar{q}q\rangle\,,$
$\langle\alpha_{s}GG\rangle$ $=$ $(6.35\pm 0.35)\times 10^{-2}\rm{GeV}^{4}\,,$ (60)
$\langle g_{s}^{3}G^{3}\rangle$ $=$ $(8.2\pm 1.0)\times\langle\alpha_{s}GG\rangle~{}\rm{GeV}^{2}\,,$
$\langle g_{s}\bar{q}\sigma Gq\rangle$ $=$ $(0.8\pm 0.2)\times\langle\bar{q}q\rangle~{}\rm{GeV}^{2}\,,$
$\langle g_{s}\bar{s}\sigma Gs\rangle$ $=$ $(0.8\pm 0.2)\times\langle\bar{s}s\rangle~{}\rm{GeV}^{2}\,.$
Additionally, we use the following values for the running masses of the charm and strange quarks at the renormalization scale of $2$ GeV Chen:2013zia :
$\alpha_{s}(\mu=M_{\tau})$ $=$ $0.33\,,$
$\alpha_{s}(\mu=2~{}{\rm GeV})$ $=$ ${\alpha_{s}(M_{\tau})\over 1+{25\alpha_{s}(M_{\tau})\over 12\pi}\log({\mu^{2}\over M_{\tau}^{2}})}=0.31\,,$
$\overline{m}_{c}(\mu=m_{c})$ $=$ ${1.23\pm 0.09}~{}\rm{GeV}\,,$ (61)
$m_{c}(\mu=2~{}{\rm GeV})$ $=$ $\overline{m}_{c}({\alpha_{s}(\mu)\over\alpha_{s}(\overline{m}_{c})})^{12/25}=1.10\,,$
$m_{s}(\mu=2~{}{\rm GeV})$ $=$ $93^{+11}_{-~{}5}~{}\rm{MeV}\,.$
The sum rule equations listed in Eq. (10), Eq. (50), and Eq. (50) contain two free parameters: the threshold value $\omega_{c}$ and the Borel mass $T$. Their appropriate working regions are determined based on three criteria: a) sufficient convergence of the OPE, b) a sufficiently large pole contribution, and c) weak dependence of the sum rule results on these parameters.
Figure 2: CVG(′,′′) and PC as functions of the Borel mass $T$, with the threshold value set to be $\omega_{c}=2.95$ GeV. These curves are derived using the current $J_{\Sigma_{cg}^{+},\bar{3}_{c},1,1,1/2^{+}}$, which belongs to the doublet $[\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
To ensure sufficient convergence of the OPE, we require the $g_{s}^{4}$ terms to be less than $5\%$, the $D=8$ terms to be less than $10\%$, and the $D=6$ terms to be less than $20\%$:
$\rm{CVG}$$\equiv$ $|{\Pi^{g_{s}^{n}=4}_{+}(\infty,T)\over\Pi_{+}(\infty,T)}|\leq 5\%\,,$
$\rm{CVG^{\prime}}$$\equiv$ $|{\Pi^{D=8}_{+}(\infty,T)\over\Pi_{+}(\infty,T)}|\leq 10\%\,,$ (62)
$\rm{CVG^{\prime\prime}}$$\equiv$ $|{\Pi^{D=6}_{+}(\infty,T)\over\Pi_{+}(\infty,T)}|\leq 20\%\,.$
As shown in Fig. 2, the lower limit of the Borel mass is determined to be $T_{min}=0.37$ GeV.
To ensure a sufficiently large pole contribution, we require the pole contribution to exceed $40\%$:
$\rm{PC}$$\equiv$ $|{\Pi_{+}(\omega_{0},T)\over\Pi_{+}(\infty,T)}|\geq 40\%\,.$ (63)
As shown in Fig. 2, the upper limit of the Borel mass is determined to be $T_{max}=0.40$ GeV, when setting $\omega_{c}=2.95~{}\rm{GeV}$.
Altogether, we obtain the Borel window $0.37~{}\rm{GeV}\leq T\leq 0.40~{}\rm{GeV}$ for $\omega_{c}=2.95~{}\rm{GeV}$. By varying $\omega_{c}$, we find that non-vanishing Borel windows exist for $\omega_{c}\geq\omega_{c}^{min}=2.7~{}\rm{GeV}$. We choose $\omega_{c}$ to be approximately $10\%$ larger and establish its working region as $2.7~{}\rm{GeV}\leq\omega_{c}\leq 3.2~{}\rm{GeV}$, where the numerical results are derived as:
$\bar{\Lambda}_{+}$ $=$ $2.46^{+0.30}_{-0.14}~{}\rm{GeV}\,,$ (64)
$f_{+}$ $=$ $0.079^{+0.033}_{-0.030}~{}\rm{GeV}^{4}\,.$ (65)
$K_{+}$ $=$ $-2.72^{+0.19}_{-0.26}~{}\rm{GeV}^{2}\,,$ (66)
$\Sigma_{+}$ $=$ $2.86^{+0.31}_{-0.25}~{}\rm{GeV}^{2}\,.$ (67)
Their variations are shown in Fig. 3 as functions of the Borel mass $T$, where the dependence is weak and remains within an acceptable range inside the Borel window $0.37~{}\rm{GeV}\leq T\leq 0.40~{}\rm{GeV}$. From this analysis, we extract the mass of the $\Sigma_{cg}^{+}$ state as
$M_{+}$ $=$ $3.46^{+0.33}_{-0.21}~{}\rm{GeV}\,,$ (68)
whose dependence on the threshold value $\omega_{c}$ and the Borel mass $T$ is illustrated in Fig. 4. The uncertainties originate from variations in the threshold value $\omega_{c}$, the Borel mass $T$, and the QCD parameters provided in Eqs. (60) and (61).
Figure 3: Variations of $\bar{\Lambda}_{+}$, $K_{+}$, and $\Sigma_{+}$ as functions of the Borel mass $T$, where the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~{}\mathrm{GeV}$, $2.95~{}\mathrm{GeV}$, and $3.05~{}\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $[\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
Figure 4: Variations of $M_{+}$ as functions of the threshold value $\omega_{c}$ and the Borel mass $T$. In the left panel the long-dashed, solid, and short-dashed curves correspond to fixed Borel masses $T=0.37~{}\mathrm{GeV}$, $0.38~{}\mathrm{GeV}$, and $0.40~{}\mathrm{GeV}$, respectively. In the right panel the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~{}\mathrm{GeV}$, $2.95~{}\mathrm{GeV}$, and $3.05~{}\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $[\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
Similarly, we apply the QCD sum rule method to study the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, $[6_{f},\bar{3}_{c},1,2,G]$, and $[6_{f},\bar{3}_{c},1,0,G]$. The obtained results are summarized in Table. 1. Additionally, we have examined the charmed hybrid baryons belonging to the other four multiplets—$[6_{f},6_{c},0,1,G]$, $[\bar{3}_{f},6_{c},1,0,G]$, $[\bar{3}_{f},6_{c},1,1,G]$, and $[\bar{3}_{f},6_{c},1,2,G]$—as well as those with the negative parity. We find that they have significantly larger masses, primarily because the relevant two-point correlation functions are negative and thus non-physical when $\omega$ is not sufficiently large.